OpenStudy (anonymous):

Find an equation of a line perpendicular to the graph of 9x–3y=7 that passes through the point at (–6, 6)

hartnn (hartnn):

first can u find the slope of line 9x-3y=7 ?

OpenStudy (anonymous):

@hartnn ya its x1 y1 x2 y2

hartnn (hartnn):

whats that??

OpenStudy (anonymous):

@hartnn thats the formula

OpenStudy (anonymous):

y=3x-7/3 is the equation in y= form

hartnn (hartnn):

no it isn't... first isolate y from that equation 9x-3y=7 and tell me what u get....

hartnn (hartnn):

right,so the slope now is co-efficient of x,which is?

OpenStudy (anonymous):

I'm not sure

OpenStudy (anonymous):

y=-1/3x+4 is my answer?

hartnn (hartnn):

the co-efficient of x is 3,so the slope is 3. yes it is :)

OpenStudy (anonymous):

you saw that its a negative 1/3 right? just wanna make sure so you dont approve the wrong answer lol thats happened before

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

thats correct,because slope of perpendicular lines are negative reciprocal

OpenStudy (anonymous):

thanks!

hartnn (hartnn):

welcome :)